# SPACE: SPike-Aware Consistency Enhancement for Test-Time Adaptation in Spiking Neural Networks

### v1

@@ 4 Experimental Results
Datasets. The experiments here were conducted on three benchmark datasets: CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C [15], each containing a variety of common image corruptions. CIFAR-10-C includes 18 corruption types, while CIFAR-100-C features 19 types, covering weather, noise, and digital distortions. Both CIFAR-10-C and CIFAR-100-C consist of 10 and 100 classes of 32$\times$32 color images, respectively, similar to CIFAR-10 and CIFAR-100 [24]. Tiny-ImageNet-C, derived from the Tiny-ImageNet dataset, contains 15 corruption types and 200 classes of 64$\times$64 color images. In addition, we also considered a neuromorphic dataset, DVS Gesture-C [18], and the related experiments are provided in the Supplementary Material Sec. 7.
Experimental Setup. All experiments were conducted under the highest corruption level (level=5). Following the experimental setup of MEMO [43], we applied AugMix [16] augmentation with a batch size of 32 to each test sample to generate a diverse set of augmented samples. The backbone model for most experiments is an SNN-based VGG architecture with Batch-Normalization Through Time (BNTT) [22], featuring BN layers at each time step, with 25 inference steps. To demonstrate the general applicability of the proposed SPACE method, additional experiments were conducted using an SNN-based ResNet-11 model [26] on the CIFAR-10-C dataset, with inference steps set to 30.
Baseline TTA Methods. We compared our approach with models pre-trained on clean datasets without any test-time adaptation (referred to as w.o. TTA). Moreover, since few TTA methods are specifically designed for SNNs, we selected two most representative traditional TTA methods, MEMO [43] and SITA [21], as our baselines.
@@ 4.1 Main Results
Shifted Domain w.o. TTA SITA MEMO SPACE
noise gaussian 72.38% 73.06% 77.73% 77.98%
shot 74.70% 74.15% 79.50% 79.34%
speckle 71.75% 71.67% 77.28% 77.38%
impulse 58.57% 58.41% 65.74% 69.41%
blur defocus 63.05% 62.94% 65.61% 71.59%
gaussian 57.03% 57.36% 59.93% 68.31%
motion 64.44% 64.36% 66.24% 72.14%
zoom 71.33% 70.72% 72.61% 74.67%
weather snow 76.32% 76.67% 77.38% 78.43%
fog 43.57% 43.40% 45.47% 52.80%
frost 75.72% 75.24% 78.64% 79.59%
digital brightness 82.44% 82.51% 83.05% 83.22%
contrast 22.54% 22.02% 23.83% 23.85%
elastic_transform 75.01% 74.74% 76.24% 75.49%
pixelate 70.25% 69.62% 73.25% 76.24%
jpeg_compression 84.28% 84.16% 86.05% 82.88%
spatter 78.42% 78.61% 79.75% 77.52%
saturate 80.85% 80.97% 82.03% 82.54%
Average 67.93% 67.81% 70.53% 72.41%
Table 1: Performance comparison of different methods on CIFAR-10-C dataset at level 5 corruption, using the VGG9 architecture.
In this section, we compare the performance of the proposed SPACE method with two existing approaches: MEMO [43] and SITA [21], which are current state-of-the-art methods for source-free and single-instance TTA. While MEMO improves TTA performance by enforcing consistency among augmented samples, it faces significant challenges in SNNs due to the loss of temporal information when relying on backpropagation. Similarly, SITA, despite not using backpropagation, fails to capture the intrinsic temporal dynamics of SNNs, rendering it ineffective. To demonstrate the broad applicability of the proposed method, we evaluate all methods on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C datasets using the VGG architecture, and on CIFAR-10-C with the ResNet-11 architecture to demonstrate the broad applicability of our proposed method.
4.1.1 Performance Comparison on VGG Model
Shifted Domain w.o. TTA SITA MEMO SPACE
noise gaussian 42.51% 43.01% 43.46% 44.71%
shot 45.40% 44.83% 45.41% 46.73%
speckle 43.35% 43.14% 43.53% 44.84%
impulse 25.50% 25.48% 26.07% 27.99%
blur defocus 42.15% 42.23% 42.47% 43.35%
gaussian 37.76% 37.88% 38.29% 39.19%
motion 41.87% 42.11% 42.55% 43.46%
zoom 46.77% 46.93% 47.21% 48.42%
glass 42.84% 43.29% 43.90% 44.47%
weather snow 44.88% 44.49% 44.74% 45.70%
fog 16.51% 16.65% 16.76% 17.62%
frost 44.12% 43.84% 44.34% 44.98%
digital brightness 49.36% 49.53% 49.95% 50.41%
contrast 5.42% 5.56% 5.62% 5.69%
elastic_transform 51.33% 50.99% 51.24% 51.45%
pixelate 52.35% 52.48% 52.82% 53.81%
jpeg_compression 58.47% 58.71% 59.29% 58.96%
spatter 47.76% 47.82% 48.63% 48.95%
saturate 46.91% 47.17% 47.17% 48.22%
Average 41.33% 41.38% 41.76% 42.58%
Table 2: Performance comparison of different methods on CIFAR-100-C dataset at level 5 corruption, using the VGG11 architecture.
Shifted Domain w.o. TTA SITA MEMO SPACE
noise gaussian 12.43% 12.71% 13.64% 16.71%
shot 15.20% 15.17% 16.45% 19.44%
impulse 8.74% 8.87% 9.51% 11.72%
defocus 7.58% 7.56% 7.51% 9.62%
blur motion 14.48% 14.57% 14.38% 16.57%
zoom 13.59% 13.94% 13.92% 15.61%
glass 6.50% 6.57% 6.60% 7.31%
weather snow 15.58% 15.72% 15.97% 18.66%
fog 5.27% 5.32% 4.91% 6.01%
frost 18.06% 18.05% 17.83% 21.02%
digital brightness 15.01% 14.82% 14.74% 17.53%
contrast 1.32% 1.38% 1.44% 1.51%
elastic_transform 20.53% 21.45% 21.00% 21.96%
pixelate 31.62% 32.20% 31.39% 31.89%
jpeg_compression 31.15% 31.51% 30.82% 31.63%
Average 14.47% 14.66% 14.67% 16.48%
Table 3: Performance comparison of different methods on Tiny-ImageNet-C dataset at level 5 corruption, using the VGG11 architecture.
Shifted Domain w.o. TTA MEMO SPACE
noise gaussian 72.88% 74.21% 75.90%
shot 74.49% 75.50% 77.81%
speckle 73.12% 74.32% 76.75%
impulse 49.59% 51.35% 56.85%
blur defocus 67.77% 68.11% 68.66%
gaussian 63.57% 64.11% 64.15%
motion 65.14% 65.37% 66.02%
zoom 73.06% 73.96% 73.82%
weather snow 78.11% 78.32% 78.74%
fog 42.84% 43.47% 45.47%
frost 73.12% 73.39% 74.04%
digital brightness 81.71% 81.82% 81.28%
contrast 12.72% 13.63% 14.94%
elastic_transform 74.88% 75.54% 76.79%
pixelate 77.08% 77.54% 80.32%
jpeg_compression 82.47% 83.24% 83.96%
spatter 78.75% 79.46% 80.58%
saturate 79.40% 79.40% 78.70%
Average 67.82% 68.49% 69.71%
Table 4: Performance comparison of different methods on CIFAR-10-C dataset at level 5 corruption, using the ResNet11 architecture.
The results on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C at level 5 corruption, shown in Tab. 1, Tab. 2, and Tab. 3, demonstrate the effectiveness of SPACE. SPACE outperforms w.o. TTA, MEMO, and SITA in average accuracy across all datasets. On CIFAR-10-C, SPACE achieves 72.41%, improving by 4.48% over w.o. TTA (67.93%), 4.60% over SITA (67.81%), and 1.88% over MEMO (70.53%). On CIFAR-100-C, SPACE reaches 42.58%, surpassing w.o. TTA (41.33%), SITA (41.38%), and MEMO (41.76%). On Tiny-ImageNet-C, SPACE achieves 16.48%, compared to 14.47% for w.o. TTA, 14.66% for SITA, and 14.67% for MEMO. Notably, SPACE consistently improves performance on Tiny-ImageNet-C across all domains, showing its robustness on complex datasets.
Across individual corruption types, SPACE consistently outperforms w.o. TTA, MEMO, and SITA. For instance, on CIFAR-10-C, SPACE shows notable improvements, such as in impulse noise (69.41% for SPACE vs. 58.57% for w.o. TTA, 58.41% for SITA, and 65.74% for MEMO) and fog (52.80% for SPACE vs. 43.57% for w.o. TTA, 43.40% for SITA, and 45.47% for MEMO). On CIFAR-100-C, SPACE similarly excels, including in motion blur (43.46% for SPACE vs. 41.87% for w.o. TTA, 42.11% for SITA, and 42.55% for MEMO).
A particularly striking result is SPACE’s performance on Tiny-ImageNet-C, where it surpasses all methods across every corruption type. For example, in gaussian noise and impulse noise, SPACE leads by a significant margin. Even in simpler corruptions like brightness and pixelate, SPACE outperforms the w.o. TTA, demonstrating its robustness in complex datasets like Tiny-ImageNet-C.
Although SPACE generally outperforms MEMO and SITA, MEMO does slightly better in a few cases, such as jpeg_compression on CIFAR-10-C (86.05% for MEMO vs. 82.88% for SPACE) and on CIFAR-100-C (59.29% for MEMO vs. 58.96% for SPACE). These minor differences suggest MEMO has an edge with specific corruptions, but they do not diminish SPACE’s overall superiority.
4.1.2 Performance Comparison on ResNet Model
We evaluated the SPACE method on the CIFAR-10-C dataset using the ResNet11 architecture, which lacks BN layers, making SITA inapplicable. The results, shown in Tab. 4, demonstrate SPACE’s effectiveness across different model structures. SPACE achieves an average accuracy of 69.71%, improving by 1.22% over MEMO (68.49%) and 1.89% over w.o. TTA (67.82%).
In terms of individual corruption types, SPACE outperforms MEMO and w.o. TTA in most cases. Notable improvements include challenging corruptions such as impulse noise (56.85% vs. 49.59% for w.o. TTA and 51.35% for MEMO) and contrast (14.94% vs. 12.72% for w.o. TTA and 13.63% for MEMO). SPACE also achieves 80.32% on pixelate, outperforming MEMO (77.54%) and w.o. TTA (77.08%). These trends align with those observed in the VGG-based experiments, reaffirming SPACE’s consistency and generalizability.
While MEMO marginally outperforms SPACE on zoom (73.96% vs. 73.82%) and brightness (81.82% vs. 81.28%), these differences are minimal and do not overshadow SPACE’s overall advantages. The results further confirm that SPACE is not dependent on a specific architecture and can adapt effectively to various backbone models.
4.2 Ablation Study
Figure 3: Local Feature Maps Based SPACE Outperforms Global Feature Maps Based SPACE.
Local vs. Global Feature Map Similarity. We compared the performance of local versus global feature map similarity for our proposed method. The results in Fig. 3 clearly illustrate the advantage of using similarity of local feature maps over global feature maps. While the global approach does not account for variability in neuron activations across channels, the local method isolates the contributions of active neurons within each channel, offering a more stable and meaningful measure of feature map consistency. As shown in Fig. 3, the local method yields a substantial accuracy improvement, whereas the global approach negatively impacts performance across all three datasets. This demonstrates that emphasizing individual channel stability leads to better generalization and adaptation. These findings underscore the distinct roles of each channel in SNNs and highlight the limitations of comparing entire feature maps.
Figure 4: Distribution of similarity values between spike counts based feature maps before and after SPACE adaptation.
Effect of SPACE on Feature Map Similarity. To further investigate the impact of SPACE adaptation on the feature map similarity, we analyzed the distribution of similarity values (computed using Eq. 4 $\sim$ Eq. 6) across the CIFAR-10-C dataset. As shown in Fig. 4, we observe a noticeable increase in the similarity between the original samples and their augmented counterparts after applying SPACE adaptation, particularly in the Gaussian Blur and Impulse Noise domains. The similarity distributions between spike counts based feature maps in these domains indicate that the feature map consistency improves following the adaptation process. Specifically, the “After SPACE Adaptation” distributions shift towards higher similarity values, reflecting the enhanced alignment of feature maps across augmented samples. This demonstrates the effectiveness of SPACE adaptation in increasing the stability of feature representations and improving generalization under various augmentation conditions.
5 Conclusion
In this work, we propose SPike-Aware Consistency Enhancement (SPACE), the first source-free and single-instance TTA method tailored for SNNs. SPACE leverages SNNs’ unique spike dynamics to optimize spike-behavior-based feature map consistency across augmented samples, addressing the limitations of ANN-based TTA methods, which often ignore the temporal and sparse nature of SNN computation. Extensive experiments on CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C, and DVS Gesture-C show SPACE achieves consistent performance improvements under severe corruptions, while generalizing effectively to various architectures like VGG9 and ResNet11. These results demonstrate SPACE’s robustness and adaptability to domain shifts, paving the way for its potential use in neuromorphic hardware and energy-efficient AI systems.
Supplementary Material
6 Kernel-Based Consistency Regularization
To investigate whether higher-dimensional feature relationships can enhance the consistency measure, we integrate kernel embeddings into the loss function. The channel-wise feature probability distribution $\mathbf{P}_{c}$ obtained from Eq. 4 can be mapped into a higher-dimensional reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, using a Gaussian kernel defined as

### v2

@@ 4 Experimental Results
Datasets and Models
The experiments were conducted on six benchmark datasets for out-of-distribution generalization: CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C [15], ImageNet-V2 [35], ImageNet-R [17], and ImageNet-A [18]. The first three datasets include 15 corruption types across four categories and five severity levels, while the three ImageNet-based datasets capture real-world domain shifts, such as distributional discrepancies and semantic variations. Additionally, we evaluated on the neuromorphic dataset DVS Gesture-C [20], which introduces six corruption types. Four backbone models were used: 1) SNN-VGG with Batch-Normalization Through Time (BNTT)[24], using 25 inference steps; 2) SNN-ResNet with 30 inference steps[27]; 3) Spike-driven Transformer V3 [44], containing 19M parameters; and 4) SNN-ConvLSTM. All models were pre-trained on their respective source domains to ensure fair evaluation on datasets with domain shifts.
Compared Methods
We compared our approach with models pre-trained on clean datasets and evaluated without any test-time adaptation (referred to as No Adapt). Additionally, since there are few TTA methods specifically designed for SNNs, we selected two representative traditional TTA methods, MEMO [47] and SITA [23], as baselines. These methods represent the current state-of-the-art for source-free and single-instance TTA, respectively. MEMO enforces output consistency among augmented samples, making it naturally compatible with SNNs by directly adjusting their final outputs. SITA, designed for models with BN layers, was adapted to SNN-VGG with BNTT by updating BN layer parameters during TTA.
Implementation details
To maximize the effectiveness of our proposed method, all experiments conducted on CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C , and DVS Gesture-C were performed with the highest severity level (level=5). In line with our setting and practical scenarios, the test batch size for all experiments was fixed to 1, and the model was updated only once per test sample. Following MEMO’s experimental setup, we applied AugMix [16] augmentation to generate diverse augmented samples, using a batch size of 32 for CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C, and DVS Gesture-C, and 64 for the ImageNet series. Please refer to Appendix for more details.
@@ 4.1 Main Results
Table 1: Performance comparison on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C with CNN-based SNN models regarding Accuracy (%). The bold number indicates the best result.
Method Noise Blur Weather Digital Average
Gauss. Shot Impl. Defoc. Glass Motion Zoom Snow Fog Frost Brit. Contr. Elas. Pix. JPEG Acc.
CIFAR-10-C, SNN-VGG9
No Adapt 72.38 74.70 58.57 63.05 63.96 64.44 71.33 76.32 43.57 75.72 82.44 22.54 75.01 70.25 84.28 66.57
SITA 73.06 74.15 58.41 62.94 64.21 64.36 70.72 76.67 43.40 75.24 82.51 22.02 74.74 69.62 84.16 66.41
MEMO 77.73 79.50 65.74 65.61 67.50 66.24 72.61 77.38 45.47 78.64 83.05 23.83 76.45 73.25 85.05 69.20
SPACE (ours) 77.98 79.34 69.41 71.59 67.76 72.14 74.67 78.43 52.80 79.59 83.22 23.85 75.49 76.24 82.88 71.03
CIFAR-10-C, SNN-ResNet11
No Adapt 72.88 74.49 49.59 67.77 63.53 65.14 73.06 78.11 42.84 73.12 81.71 12.72 74.88 77.08 82.47 65.96
MEMO 74.21 75.50 51.35 68.11 64.65 65.37 73.96 78.32 43.47 73.39 81.82 13.63 75.54 77.54 83.24 66.67
SPACE (ours) 75.90 77.81 56.85 68.66 67.64 66.02 73.82 78.74 45.47 74.04 81.28 14.94 76.79 80.32 83.96 68.15
CIFAR-100-C, SNN-VGG11
No Adapt 42.51 45.40 25.50 42.15 42.84 41.87 46.77 44.88 16.51 44.12 49.36 5.42 51.33 52.35 58.47 40.63
SITA 43.01 44.83 25.48 42.23 43.29 42.11 46.93 44.49 16.65 43.84 49.53 5.56 50.99 52.48 58.71 40.68
MEMO 43.46 45.41 26.07 42.47 43.90 42.55 47.21 44.74 16.76 44.34 49.95 5.62 51.24 52.82 59.29 41.06
SPACE (ours) 44.71 46.73 27.99 43.35 44.47 43.46 48.42 45.70 17.62 44.98 50.41 5.69 51.45 53.81 58.96 41.58
Tiny-ImageNet-C, SNN-VGG11
No Adapt 12.43 15.20 8.74 7.58 6.50 14.48 13.59 15.58 5.27 18.06 15.01 1.32 20.53 31.62 31.15 14.47
SITA 12.71 15.17 8.87 7.56 6.57 14.57 13.94 15.72 5.32 18.05 14.82 1.38 21.45 32.20 31.51 14.66
MEMO 13.64 16.45 9.51 7.51 6.60 14.38 13.92 15.97 4.91 17.83 14.74 1.44 21.00 31.39 30.82 14.67
SPACE (ours) 16.71 19.44 11.72 9.62 7.31 16.57 15.61 18.66 6.01 21.02 17.53 1.51 21.96 31.89 31.63 16.48
In this section, we compare the performance of the proposed SPACE method with two existing approaches: MEMO and SITA. To demonstrate the broad applicability of the proposed method, we evaluate all methods on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C datasets using SNN-VGG, on CIFAR-10-C with SNN-ResNet11, on ImageNet series with Spike-driven Transformer V3, and on DVS Gesture-C with SNN-ConvLSTM, to demonstrate the broad applicability of our proposed method.
Performance Comparison on CNN-based SNN Models
Figure 3: Performance comparison on ImageNet-V2/R/A with Spike-driven Transformer V3 regarding Accuracy (%). The bold number indicates the best result.
Method Accuracy (%)
V2 R A Avg.
No Adapt 67.64 41.94 15.12 41.57
MEMO 67.57 42.15 14.97 41.56
SPACE (ours) 68.82 45.07 16.22 43.37
The results on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C at level 5 corruption, shown in Table 1, highlight the effectiveness of SPACE. Key observations include the following: 1) Superior Performance Across Methods: SPACE consistently outperforms No Adapt, SITA, and MEMO in average accuracy across all datasets under both VGG and ResNet architectures. 2) Broad Applicability: SPACE demonstrates effectiveness across different CNN structures, showcasing its generalizability. 3) Robustness on Complex Datasets: A particularly notable result is SPACE’s performance on Tiny-ImageNet-C, where it surpasses all methods across every corruption type. This highlights its robustness in handling complex datasets like Tiny-ImageNet-C. 4) Limitations of SITA on SNNs: The SNN-ResNet does not include BN layers, rendering SITA inapplicable. Furthermore, SITA performs worse than No Adapt, indicating its ineffectiveness for SNNs. This may be attributed to the spike-based transmission in SNNs: directly adjusting BN parameters without gradient guidance can disrupt spike information, leading to significant performance degradation. 5) Comparison with MEMO: While SPACE generally outperforms MEMO and SITA, MEMO exhibits slightly better performance in a few specific cases, such as JPEG corruption on CIFAR-10-C (85.05% for MEMO vs. 82.88% for SPACE) and CIFAR-100-C (59.29% for MEMO vs. 58.96% for SPACE). These minor differences suggest MEMO may have an advantage under certain corruption types, but they do not detract from SPACE’s overall superiority.
Performance Comparison on Spike-driven Transformer V3
To explore the real-world impact of SPACE, we conducted tests on diverse datasets, including ImageNet-V2, ImageNet-R, and ImageNet-A. Furthermore, to evaluate its adaptability in deeper SNN architectures, we tested SPACE on the state-of-the-art Spike-Event Transformer V3. As shown in Figure 3, SPACE consistently outperforms MEMO, achieving a 1.80% improvement, which demonstrates its robustness in real-world scenarios and its effectiveness in advanced SNN models.
Performance Comparison on ConvLSTM-based SNN Model
SNNs are inherently well-suited for processing event-based neuromorphic data, thanks to their temporal dynamics and energy efficiency. The DVS-Gesture dataset [1] provides an ideal benchmark to evaluate the performance of SPACE on neuromorphic datasets, highlighting its compatibility with dynamic vision tasks and data with temporal structures. To further explore potential network architectures, we incorporated ConvLSTM. As shown in Table 2, our approach achieves significantly better results compared to existing methods, demonstrating its effectiveness and adaptability.
Table 2: Performance comparison on DVS Gesture-C with ConvLSTM-based SNN model regarding Accuracy (%). The bold number indicates the best result.
Method Accuracy (%)
DropPixel DropEvent RefractoryPeriod TimeJitter SpatialJitter UniformNoise Avg.
No Adapt 53.03 54.55 36.74 31.06 92.05 91.67 59.85
MEMO 56.06 55.30 36.74 30.68 90.91 91.67 60.23
SPACE (ours) 57.20 57.95 52.27 43.18 92.42 92.42 65.91
4.2 Ablation Studies
(a) Local feature maps based SPACE outperforms global feature maps based SPACE.
(b) Similarity distribution of spike count feature maps before and after SPACE adaptation.
Figure 4: Results of ablation study.
Local vs. Global Feature Map Similarity
We compared the performance of local versus global feature map similarity for our proposed method. The results in 4(a) clearly illustrate the advantage of using similarity of local feature maps over global feature maps. While the global approach does not account for variability in neuron activations across channels, the local method isolates the contributions of active neurons within each channel, offering a more stable and meaningful measure of feature map consistency. As shown in 4(a), the local method yields a substantial accuracy improvement, whereas the global approach negatively impacts performance across all three datasets. This demonstrates that emphasizing individual channel stability leads to better generalization and adaptation. These findings underscore the distinct roles of each channel in SNNs and highlight the limitations of comparing entire feature maps.
Effect of SPACE on Feature Map Similarity
To further investigate the impact of SPACE adaptation on the feature map similarity, we analyzed the distribution of similarity values (computed using Equation 4) across the CIFAR-10-C dataset. As shown in 4(b), we observe a noticeable increase in the similarity between the original samples and their augmented counterparts after applying SPACE adaptation, particularly in the Impulse Noise domain. The similarity distributions between spike counts based feature maps in these domains indicate that the feature map consistency improves following the adaptation process. Specifically, the "After SPACE Adaptation" distributions shift towards higher similarity values, reflecting the enhanced alignment of feature maps across augmented samples. This demonstrates the effectiveness of SPACE adaptation in increasing the stability of feature representations and improving generalization under various augmentation conditions.
5 Conclusion and Limitation
In this work, we propose SPike-Aware Consistency Enhancement (SPACE), the first source-free and single-instance TTA method tailored for SNNs. By leveraging the unique spike-driven dynamics of SNNs, SPACE optimizes spike-behavior-based feature map consistency across augmented samples, addressing the limitations of ANN-based TTA methods that overlook the temporal and sparse nature of SNN computations. Experiments on CIFAR-10, CIFAR-100, Tiny-ImageNet, ImageNet, and DVS Gesture series datasets show that SPACE consistently improves performance under severe corruptions and generalizes well across architectures, demonstrating robustness to domain shifts. However, SPACE’s reliance on backpropagation may limit its applicability in resource-constrained scenarios. Future work will explore alternative optimization strategies to expand its practicality for neuromorphic hardware and energy-efficient AI.
Appendix A Kernel-Based Consistency Regularization
To investigate whether higher-dimensional feature relationships can enhance the consistency measure, we integrate kernel embeddings into the loss function. The channel-wise feature probability distribution $\mathbf{P}_{c}(\mathbf{x}_{i})$ can be mapped into a higher-dimensional reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, using a Gaussian kernel defined as